OpenStudy (henrietepurina):

Algebra 1 question :D Medal + Fan!

OpenStudy (henrietepurina):

A basketball is thrown upwards. The height f(t), in feet, of the basketball at time t, in seconds, is given by the following function: f(t) = -16t^2 + 16t + 32 Which of the following is a reasonable domain of the graph of the function when the basketball falls from its maximum height to the ground? 0.5 < t < 1 0.5 < t < 2 1 < t < 2 1 < t < 1.5

OpenStudy (henrietepurina):

@jagr2713

jagr2713 (jagr2713):

highest point in this case should be the vertex and the rightest zero x value of vertex = -b/2a given ax^2 + bx + c

jagr2713 (jagr2713):

plug in and that'll be the first point, the other point would be solve -16x^2 + 16x + 32 = 0 and take the positive answer, that'll be your other point

OpenStudy (henrietepurina):

ok ill try...

OpenStudy (henrietepurina):

the 2nd point is 2...

OpenStudy (henrietepurina):

can you help me get the first one?

jagr2713 (jagr2713):

so if its 2 then what can we eliminate

OpenStudy (henrietepurina):

a and d

jagr2713 (jagr2713):

solve this -16x^2 + 16x + 32 = 0

jagr2713 (jagr2713):

first factor out the greatest common factor (-16) then factor... do you know how to do that?

OpenStudy (henrietepurina):

somewhat... but I'll need help

jagr2713 (jagr2713):

you will get 0=-16(t+0.5)+6

OpenStudy (henrietepurina):

ok, and by using my amazing brain smarts, I see a 0.5 in there

OpenStudy (henrietepurina):

:D

OpenStudy (henrietepurina):

making me think it might be b?

jagr2713 (jagr2713):

Yes :D

OpenStudy (henrietepurina):

thanks :D

jagr2713 (jagr2713):

yw

OpenStudy (funnyguy2):

is this a test

OpenStudy (henrietepurina):

why?

OpenStudy (funnyguy2):

just tell me first

OpenStudy (henrietepurina):

no, my algebra 1 homework....

OpenStudy (funnyguy2):

ok then your fine

OpenStudy (henrietepurina):

why would you ask that?

jagr2713 (jagr2713):

why ?@funnyguy2

OpenStudy (henrietepurina):

he/she left :|

OpenStudy (funnyguy2):

your not allowed to get help on a test

jagr2713 (jagr2713):

??? where does it say that

OpenStudy (henrietepurina):

mhm... cool, but as you know... this is NOT a test

OpenStudy (henrietepurina):

and yeah... where does it say that?

OpenStudy (henrietepurina):

hello @funnyguy2 ?

OpenStudy (funnyguy2):

@pooja195 are you allowed to help someone on a test

OpenStudy (funnyguy2):

@sleepyjess

OpenStudy (funnyguy2):

@TheSmartOne

OpenStudy (sleepyjess):

It is not a test, they already said that.

OpenStudy (henrietepurina):

exactly....

OpenStudy (funnyguy2):

one question per post please

OpenStudy (henrietepurina):

but what he asked was, are you allowed to help IF it was a test?

OpenStudy (henrietepurina):

are you kidding @funnyguy2 ?

pooja195 (pooja195):

@funnyguy2 it is certainly not allowed unless youve made a mistake by accidently halping them without knowing its a test i am guilty of that but if you have helped knowing its a test you can land in some serious trouble

OpenStudy (henrietepurina):

ok thanks for clearing that up :D

jagr2713 (jagr2713):

close question pls :D

OpenStudy (henrietepurina):

sure can do :D

TheSmartOne (thesmartone):

You guys asked where that is? :P Let me pull it up :) Even though I know you weren't cheating :)

TheSmartOne (thesmartone):

The term and conditions are something you agreed to when you first signed up! And in the TaC (Terms and Conditions) it has near the bottom: http://prntscr.com/5wy0p9 Don't want to copy-paste otherwise it looks longer than it is... lol Anyways here is the whole TaC: http://openstudy.com/terms-and-conditions